OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 7 times x minus 21 all over 18 minus 6 times x.? [Hint: (a−x) = − (x − a)]

OpenStudy (anonymous):

These are the multiple choice answers 6 over 7. x − 3 7 over 6. x + 3

hero (hero):

\[\frac{7x - 21}{18 - 6x} = \frac{7(x - 3)}{6(3 - x)} =\frac{7(x-3)}{-6(x - 3)} \]

OpenStudy (anonymous):

how did you get -6(x-3)

hero (hero):

6(3 - x) = -6(x - 3) Remember the Hint you were given above?

hero (hero):

What expression cancels?

OpenStudy (anonymous):

would the answer be 7 overs -6

hero (hero):

Which expression cancels? Do you see a factor of one that cancels?

hero (hero):

Do you know what it means to cancel a factor of one?

OpenStudy (anonymous):